# Conformalized Ordinal Classification with Marginal and Conditional Coverage

### v1

@@ 1 Introduction
Conformal prediction is a unique distribution-free UQ technique that provides a prediction set rather than a point prediction for the true response with guaranteed coverage (Vovk et al., 1999, 2005; Shafer and Vovk, 2008; Angelopoulos and Bates, 2021; Fontana et al., 2023). It can be used as a wrapper with any black-box algorithm. In this paper, we use the conformal prediction technique to construct prediction sets for ordinal classification problems. By combining the ideas of conformal prediction and multiple testing, two new conformal prediction methods are introduced for constructing contiguous and non-contiguous prediction sets. Firstly, the problem of ordinal classification is reformulated as a problem of multiple testing; Secondly, for each constructed hypothesis, the marginal and conditional conformal $p$-values are respectively calculated; Thirdly, based on these marginal (conditional) conformal $p$-values, three multiple testing procedures are developed for controlling marginal (conditional) familywise error rate (FWER); Finally, based on the testing outcomes of these procedures, the prediction sets are constructed and proved having guaranteed marginal (conditional) coverage.
There are almost no works of applying conformal prediction to address ordinal classification in the literature. To our knowledge, Lu et al. (2022) is the only existing work, in which, a new (split) conformal prediction method is developed for constructing adaptive contiguous prediction region. This method is proved to have guaranteed marginal coverage, however, it cannot guarantee to have more desired conditional coverage. Moreover, it does not work well for high dimensional data. Compared to the method introduced in Lu et al. (2022), our proposed methods generally show via theoretical and numerical studies performing better in the settings of higher dimensions and in terms of class-specific conditional coverage; especially for the conditional conformal $p$-values based methods, they are proved to have guaranteed conditional coverage.
The rest of this paper is structured as follows. In Section 2, we briefly introduce split-conformal prediction and review related works, followed by Section 3 which presents the development of our proposed conformal methods using the idea of multiple testing. Section 4 provides numerical studies to evaluate the performance of the proposed methods compared to the existing method. Some discussions are presented in Section 5 and all proofs
@@ 2.1 Conformal Prediction
Conformal prediction is a general approach to construct prediction sets combined with a pre-trained classifier. The main advantage of this approach is that it is distribution-free and can work with any black-box algorithm. Conformal prediction is broadly of two types – full conformal prediction and split-conformal prediction. The full conformal prediction uses all the observations to train the black-box algorithms (Vovk et al., 2005). In contrast, split-conformal prediction (Papadopoulos et al., 2002) involves splitting the training data into proper training data to train the black-box algorithm and calibration data to calculate the threshold for forming prediction sets. Our proposed methods are based on the split-conformal method. Consider a multi-class classification problem with feature space $\mathcal{X}$ and labels $\mathcal{Y}=\{1,2,\cdots,K\}$. Given the training observations $(X_{i},Y_{i})_{i=1}^{2n}$ and a test input $X_{2n+1}$, the goal is to find a prediction set $C(X_{2n+1}):\mathcal{X}\mapsto 2^{\mathcal{Y}}$ that contains the unknown response $Y_{2n+1}$ with enough statistical coverage.
The split-conformal procedure suggests to split $2n$ observations to $n$ training observations, i.e., $(X_{i},Y_{i})_{i=1}^{n}$, which are used to train $\hat{f}$, a black-box classifier such that $\hat{f}:\mathcal{X}\mapsto\mathcal{Y}$ and the remaining $n$ observations $(X_{i},Y_{i})_{i=n+1}^{2n}$ for calibration. The central part of this technique involves calculating the conformity scores for each observation, which measures how much the test observation conforms with the calibration observations. There can be several choices of conformity scores for multi-class classification problem, including posterior class probability, cumulative probability, and regularized cumulative probability (Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020). Given the score function $s:\mathcal{X}\times\mathcal{Y}\mapsto\mathbb{R}$, the conformity score for the $i^{th}$ calibration observation is defined as $s_{i}=s(X_{i},Y_{i}),\hskip 5.69046pti=n+1,\cdots,2n$.
@@
In general, the concept of conditional coverage such as object conditional validity and class-specific conditional validity are more relevant to practical applications (Vovk, 2012; Lei, 2014; Barber et al., 2021). If it is satisfied, the more desired results are often guaranteed. Specifically, in classification problems, class-specific conditional validity provides conditional coverage for each given class, which is defined as
@@
Moreover, if the conformity scores $\{s_{i}\}_{i=n+1}^{2n+1}$ are distinct surely, we have $p(Z_{2n+1})\sim U\Bigl{\{}\frac{1}{n+1},\cdots,1\Bigl{\}}$.
@@ 2.2 Related work
The framework of Conformal prediction was introduced by Vladimir Vovk and his collaborators Vovk et al. (1999, 2005) and has found many applications in classification problems. Shafer and Vovk (2008) and Angelopoulos and Bates (2021) provided a tutorial introduction and brief literature review on this field. Several conformal methods have been developed to address binary classification (Lei, 2014) and multi-class classification problems (Hechtlinger et al., 2018; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Tyagi and Guo, 2023). Coverage guarantees of all these methods are established under the assumption of exchangeability. Very recently, some new conformal prediction methods have been developed in the settings of non-exchangeability (Tibshirani et al., 2019; Cauchois et al., 2021; Gibbs and Candes, 2021).
@@ 4.2 Application to real data
We also evaluate the performance of our proposed methods on a real dataset, Traffic accident data, which is publicly available on the website of Leeds City Council. The real data consists of 1,908 traffic accidents that occurred in the year 2019. The objective is to predict the severity of the casualties, which are classified into three categories – mild, serious, and fatal based on the features available. In the numerical experiment, 500 observations are used to train the logistic regression model, $35\%$ of the remaining observations are used for calibration, and $65\%$ for validation. Figure 4 shows that all these methods empirically achieve desired marginal coverage for different levels of mis-coverage, however, the proposed marginal OPI and marginal OPS, and existing OAPS have lower set sizes than the conditional OPI and conditional OPS. It is also evident from Figure 4 that the proposed conditional OPS and conditional OPI both attain desired class-specific conditional coverage unlike the existing method, OAPS, which seems to largely deviate from the desired level of conditional coverage.

### v2

# Distribution-free Conformal Prediction for Ordinal Classification

@@ 1 Introduction
Conformal prediction is a unique distribution-free UQ technique that provides a prediction set rather than a point prediction for the true response with guaranteed coverage (Vovk et al., 1999, 2005; Shafer and Vovk, 2008; Angelopoulos and Bates, 2021; Fontana et al., 2023). It can be used as a wrapper with any underlying algorithm. In this paper, we use the conformal prediction technique to construct prediction sets for ordinal classification problems. By combining the ideas of conformal prediction and multiple testing, two new conformal prediction methods are introduced for constructing contiguous and non-contiguous prediction sets. Firstly, the problem of ordinal classification is reformulated as a problem of multiple testing; Secondly, for each constructed hypothesis, the marginal and conditional conformal $p$-values are respectively calculated; Thirdly, based on these marginal (conditional) conformal $p$-values, three multiple testing procedures are developed for controlling marginal (conditional) familywise error rate (FWER); Finally, based on the testing outcomes of these procedures, the prediction sets are constructed and proved having guaranteed marginal (conditional) coverage. There are almost no works of applying conformal prediction to address ordinal classification in the literature. To our knowledge, Lu et al. (2022) is the only existing work, in which, a new (split) conformal prediction method is developed for constructing adaptive contiguous prediction region. This method is proved to have guaranteed marginal coverage, however, it cannot guarantee to have more desired conditional coverage. Moreover, it does not work well for high dimensional data. Compared to the method introduced in Lu et al. (2022), our proposed methods generally show via theoretical and numerical studies better performance in the settings of higher dimensions and in terms of class-specific conditional coverage; especially for the conditional conformal $p$-values based methods, they are proved to have guaranteed conditional coverage.
The rest of this paper is structured as follows. In Section 2, we briefly introduce split-conformal prediction and review related works, followed by Section 3 which presents the development of our proposed conformal methods using the idea of multiple testing. Section 4 provides numerical studies to evaluate the performance of the proposed methods compared to the existing method. Some discussions are presented in Section 5 and all proofs are deferred to the appendix section.
@@ 2.1 Conformal Prediction
Conformal prediction is a general statistical technique to construct prediction sets for any supervised learning method. The main advantage of this approach is that it is distribution-free and can work with any underlying algorithm. Conformal prediction is broadly of two types – full conformal prediction and split-conformal prediction. The full conformal prediction uses all the observations to train the underlying algorithms (Vovk et al., 2005). In contrast, split-conformal prediction (Papadopoulos et al., 2002) involves splitting the training data into proper training data to train the underlying algorithm and calibration data to calculate the threshold for forming prediction sets. Our proposed methods are based on the split-conformal method. Consider a multi-class classification problem with feature space $\mathcal{X}$ and labels $\mathcal{Y}=\{1,2,\cdots,K\}$. Given the training observations $(X_{i},Y_{i})_{i=1}^{2n}$ and a test input $X_{2n+1}$, the goal is to find a prediction set $C(X_{2n+1}):\mathcal{X}\mapsto 2^{\mathcal{Y}}$ that contains the unknown response $Y_{2n+1}$ with enough statistical coverage.
The split-conformal procedure suggests to split $2n$ observations to $n$ training observations, i.e., $(X_{i},Y_{i})_{i=1}^{n}$, which are used to train $\hat{f}$, an underlying classifier such that $\hat{f}:\mathcal{X}\mapsto\mathcal{Y}$ and the remaining $n$ observations $(X_{i},Y_{i})_{i=n+1}^{2n}$ for calibration. The central part of this technique involves calculating the conformity scores for each observation, which measures how much the test observation conforms with the calibration observations. There can be several choices of conformity scores for multi-class classification problem, including posterior class probability, cumulative probability, and regularized cumulative probability (Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020). Given the score function $s:\mathcal{X}\times\mathcal{Y}\mapsto\mathbb{R}$, the conformity score for the $i^{th}$ calibration observation is defined as $s_{i}=s(X_{i},Y_{i}),\hskip 5.69046pti=n+1,\cdots,2n$.
@@
In general, the concept of conditional coverage such as object conditional validity and class-specific conditional validity are more relevant to practical applications (Vovk, 2012; Lei, 2014; Barber et al., 2021). If conditional coverage is guaranteed, our predictions are valid for a more specific sub-population, not merely on average over a general population. Specifically, in classification problems, class-specific conditional validity provides conditional coverage for each given class, which is defined as
@@
Moreover, if the conformity scores $\{s_{i}\}_{i=n+1}^{2n+1}$ are distinct surely, we have,
$p(Z_{2n+1})\sim U\Bigl{\{}\frac{1}{n+1},\cdots,1\Bigl{\}}$
@@ 2.2 Related work
The framework of Conformal prediction was introduced by Vladimir Vovk and his collaborators Vovk et al. (1999, 2005) and has found many applications in classification problems. Shafer and Vovk (2008) and Angelopoulos and Bates (2021) provided a tutorial introduction and brief literature review on this field. Several conformal methods have been developed to address binary classification (Lei, 2014) and multi-class classification problems (Hechtlinger et al., 2018; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Tyagi and Guo, 2023). Coverage guarantees of all these methods are established under the assumption of exchangeability. Very recently, some new conformal prediction methods have been developed in the settings of non-exchangeability (Tibshirani et al., 2019; Gibbs and Candes, 2021; Barber et al., 2023).
@@ 4.2 Application to real data
We also evaluate the performance of our proposed methods on a real dataset, Traffic accident data, which is publicly available on the website of Leeds City Council [^1] The traffic accident dataset is taken from Leeds Council and its web link url is https://www.leeds.gov.uk/parking-roads-and-travel/connecting-leeds-and-transforming-travel/road-safety/road-traffic-collision-statistics.. The real data consists of 1,908 traffic accidents that occurred in the year 2019. The objective is to predict the severity of the casualties, which are classified into three categories – mild, serious, and fatal based on the features available. In the numerical experiment, 500 observations are used to train the logistic regression model, $35\%$ of the remaining observations are used for calibration, and $65\%$ for validation. Figure 4 shows that all these methods empirically achieve desired marginal coverage for different levels of mis-coverage, however, the proposed marginal OPI and marginal OPS, and existing OAPS have lower set sizes than the conditional OPI and conditional OPS. It is also evident from Figure 4 that the proposed conditional OPS and conditional OPI both attain desired class-specific conditional coverage unlike the existing method, OAPS, which seems to largely deviate from the desired level of conditional coverage.